OpenStudy (anonymous):

Anyone please? Solve the following system of equations using the elimination method. 4x+5y =-1 2x+3y=7

OpenStudy (mertsj):

Multiply the second equation by -2 and add the result to the first equation.

OpenStudy (mertsj):

That will eliminate x

OpenStudy (mertsj):

Tell me what you get.

OpenStudy (anonymous):

umm 4x-6y=-14 ?

OpenStudy (mertsj):

Just about Watch this.

OpenStudy (mertsj):

\[-2(2x+3y=7)=-4x-6y=-14\]

OpenStudy (mertsj):

Now add that to the first equation.

OpenStudy (mertsj):